OpenStudy (anonymous):

If f(x)=x^2+4x, find the value(s) of x if the f(x) is 5.

OpenStudy (anonymous):

The answer from the back of my text book is 1,-5.. how?

OpenStudy (anonymous):

I thought I have to do: \[f(x) = x ^{2}+4x\] \[5=x ^{2}\]

OpenStudy (anonymous):

+4x

OpenStudy (anonymous):

and from there on, I am stuck.

OpenStudy (anonymous):

it does for my other questions. But for this one I don't think that's how I should solve my question D:

OpenStudy (anonymous):

It's the only way :X you just substitute X by the value given, wich in this case is 5

OpenStudy (anonymous):

f(x) = x^2 + 4x f(x) = 5 when you plug it in... 5 = x^2 + 4x can you get the two values for x from here?

OpenStudy (anonymous):

5= x^2 + 4x( subtract 5 ) x^2 + 4x - 5 = 0 ( quadratic) ( x +5) ( x- 1)= 0 ( factors) x= -5 and x= 1

OpenStudy (anonymous):

no I can't because I forgot how to do this D:

OpenStudy (anonymous):

ohhhhh!!!

OpenStudy (anonymous):

factored form! thank you!!